OpenStudy (anonymous):

What logarithmic function represents the data in the table? x f(X) 25 2 125 3 625 4

OpenStudy (snowsurf):

Hint. You will need to use the relation between power and logarithms

OpenStudy (zale101):

x f(X) 25 2 what is the number, that when squared (\(number)^2\) = \(25\).

OpenStudy (zale101):

if you know that number, then you can apply it with the rest of the information given. x f(X) 25 2 \((number)^2=25\) 125 3 \((number)^3=125\) 625 4 \((number)^4=625\)

OpenStudy (zale101):

The function, therefore, would be \(f(x)=(number)^x\) that would be the rule that will apply for all the plotted points given.